# Integration of Cobalt Ferromagnetic Control Gates for Electrical and Magnetic Manipulation of Semiconductor Quantum Dots

### v1

@@ Abstract
The rise of electron spin qubit architectures for quantum computing processors has led to a strong interest in designing and integrating ferromagnets to induce stray magnetic fields for electron dipole spin resonance (EDSR). The integration of nanomagnets imposes however strict layout and processing constraints, challenging the arrangement of different gating layers and the control of neighboring qubit frequencies. This work reports a successful integration of nano-sized cobalt control gates into a multi-gate FD-SOI nanowire with nanometer-scale dot-to-magnet pitch, simultaneously exploiting electrical and ferromagnetic properties of the gate stack at nanoscale. The electrical characterization of the multi-gate nanowire exhibits full field effect functionality of all ferromagnetic gates from room temperature to 10 mK, proving quantum dot formation when ferromagnets are operated as barrier gates. The front-end-of-line (FEOL) compatible gate-first integration of cobalt is examined by energy dispersive X-ray spectroscopy and high/low frequency capacitance characterization, confirming the quality of interfaces and control over material diffusion. Insights into the magnetic properties of thin films and patterned control-gates are provided by vibrating sample magnetometry and electron holography measurements. Micromagnetic simulations anticipate that this structure fulfills the requirements for EDSR driving for magnetic fields higher than 1 T, where a homogeneous magnetization along the hard magnetic axis of the Co gates is expected. The FDSOI architecture showcased in this study provides a scalable alternative to micromagnets deposited in the back-end-of-line (BEOL) and middle-of-line (MOL) processes, while bringing technological insights for the FEOL-compatible integration of Co nanostructures in spin qubit devices.
@@
ferromagnet, cobalt, CMOS integration, quantum dots, EDSR, spin qubitsFabio Bersano* Michele Aldeghi Niccolò Martinolli Victor Boureau Thibault Aboud Michele Ghini Pasquale Scarlino Gian Salis Adrian M. Ionescu*
1Nanoelectronic Devices Laboratory (Nanolab), IEM EPFL, 1015 Lausanne, Switzerland
2IBM Research-Zurich, 8803 Rüschlikon, Switzerland
3Interdisciplinary Center for Electron Microscopy, EPFL, 1015 Lausanne, Switzerland.
4Hybrid Quantum Circuits Laboratory (HQC Lab), IPHYS EPFL, 1015 Lausanne, Switzerland
* Corresponding authors. Email addresses: fabio.bersano@epfl.ch, adrian.ionescu@epfl.ch
@@ 1 Introduction
Particularly for electron spin qubits in Si, the need for plunger, depletion/barrier gates and micromagnet integration puts stringent restrictions on layout level when scaling up the quantum processor, increasing the complexity of gating layers for electrostatic control. This challenge can be addressed by properly designing the device architecture and combining the functionalities of different metallization layers into one. While a 2D architecture proves more fitting for expansive networks [28], it necessitates additional depletion gates to induce 1D confinement, a role that is naturally fulfilled by etched structures like FinFETs and nanowires, which typically require a minimum of three metal layers [25, 10]. In this work, we propose a device architecture aimed to solve the scalability issue of EDSR-driven electron spin qubits while anticipating high addressability of neighbouring qubits. In our structure, the ferromagnetic property of cobalt nano-gates is combined with a DC (and possibly RF) field-effect control of the semiconductor potential, with 1D confinement provided by ultra-thin fully depleted silicon-on-insulator (FDSOI) nanowires (Figure 1a). The incorporation of the ferromagnetic gates suggested here is first analyzed with an emphasis on its compatibility with FEOL and middle-of-line (MOL) CMOS processes. The fabrication and electrical characterization of a proof-of-concept quantum dot device (Figure 1b) is reported, and micromagnetic simulations are discussed to anticipate the experimental conditions required for EDSR driving induced by Co nanogates.
@@ 2.1 Thin Film Morphology and Thermal Budget
In the context of semiconductor devices manufacturing, cobalt is considered a CMOS-compatible material. It has been explored as a replacement for traditional interconnects metals like tungsten and copper due to its low electron mean free path [29], and as a valid alternative to TiSi2 and NiSi to form low-resistivity silicide contacts (CoSi2) in advanced technology nodes [30]. Co interconnects have been successfully integrated for CMOS nodes below 100 nm in the MOL metallization layers and as cladding coating at higher metal levels, showing more than two times reduction in via resistance [31] and 5-10 times improvement in electromigration [32]. However, instances of extrinsic time-dependent dielectric breakdown and atomic diffusion have been documented, prompting the need for diffusion barriers [33, 34]. Even a partial diffusion of Co atoms into the gate oxide and silicon could result in the creation of interfacial silicides [35], which may lead to the development of local magnetic moments with random orientations compared to the magnetization of the thin film. Yet, it has been demonstrated that a thin SiO2 layer effectively inhibits the formation of interfacial CoSi/CoSi2 due to the strong binding of Si atoms to the thermal SiO2 layer [36]. However, aluminum oxide does not exhibit this diffusion barrier effect [33], necessitating the use of an additional adhesion layer to separate the Co from the gate oxide. Because of the partial diffusion of Cr into high-k dielectrics and reaction with oxygen to form Cr2O3, a Cr-Co alloy provides a reliable self-forming diffusion barrier that shows a breakdown voltage two times higher than pure Co [37]. Motivated by this, we first deposit a thin Cr layer (2-3 nm) before Co e-beam evaporation; this layer also enhances metal adhesion to the gate oxide and improves the lift-off process. The cross-section of a Cr/Co barrier gate and Ti/Pd plunger gate annealed for 10 minutes at 300 °C in forming gas (5 H2, 95 N2) are shown in Figure 2a, where energy dispersive X-ray (EDX) TEM spectroscopy was used to highlight the materials. No Cr/Co or Ti/Pd diffusion was observed in the oxide, indicating effective atomic screening by the thin adhesion metal, though partial diffusion of the adhesion layers into Co and Pd occurred.
In addition to inter-layer diffusion, the oxidation of Co must be considered since it occurs rapidly when in contact with oxygen [38]. Three oxide growth processes have been documented: an immediate hydroxide formation (Co(OH)2) of a few Ångstrom ( nm) upon air exposure of the surface at annealing temperatures below 100 °C, a logarithmic growth of CoO for annealing temperatures between 100-225 °C, and a quadratic growth rate of a combination of CoO and Co3O4 for temperatures exceeding 300 °C [39, 40]. We analyzed the oxidation phase of Co through angle-resolved X-ray photoelectron spectroscopy (ARXPS) of thin films (33 nm) deposited on both SiO2 and Al2O3 before and after annealing. Results are independent of the dielectric used and are reported in Figure 2b. Upon thermal treatment at 300 °C, two peaks separated by eV are measured, identifying the binding energy profile of Co and its oxides, in accordance with data reported in literature [41]. The measurement at lower energy identifies two peaks corresponding to Cr2O3, proving experimentally the diffusion of Cr into Co. This is reflected in an increase in coercivity, as confirmed by room temperature vibrating sample magnetometry (VSM) measurements in Figure 2c. It is well known that CoCr layers exhibit varying coercivity and saturation magnetization with changes in Cr concentration, which is typically increased up to 20 for recording media [42]. However, due to the low concentration, no significant change in magnetization saturation was measured in this case.
The evolution of crystallite size and crystal phase of thin Co evaporated films was studied by bright-field TEM and selected area electron diffraction (SAED) measurements [43, 44]. The results, shown in Figure 2d, indicate a grain size ranging from 15 to 20 nm with intensity peaks corresponding to the hexagonal close packed (hcp) phase, as confirmed by grazing incidence XRD (GIXRD) measurements. After annealing, a slight contribution of crystallites from the face centered cubic (fcc) phase was detected, in agreement with the increase in magnetic coercivity and with what has been reported in the literature for evaporated Co thin films [45]. Under all thermal treatments studied in this work, the surface oxidation was limited to nm by cooling the sample to room temperature in nitrogen atmosphere before exposing it to ambient conditions. Although not detrimental for the field-effect operation of the ferromagnetic gates, their partial oxidation should be taken into account for magnetic purposes, as the bulk fcc phase of CoO presents antiferromagnetic properties. Our study therefore suggests that the thermal budget for an FEOL-compatible process for the integration of Co nanomagnets should be kept below 300 °C to avoid magnetic inhomogeneities induced by oxidation and interlayer diffusion.
@@ 2.2 Work Function and Interface Traps
The primary criteria for incorporating a metal as gating layer include its chemical stability with a given dielectric to regulate the threshold voltage (Vth), negligible C-V hysteresis, and low interface traps. All these properties depend on the thermal history of the fabrication process and set boundaries to the thermal budget. We analyzed the density of traps in metal-oxide-semiconductor (MOS) capacitors fabricated with Cr-Co (3-33 nm) and Ti-Pd (3-33 nm) on thermal SiO2 (grown at 850 °C in dichloroethene atmosphere) and Al2O3 deposited by atomic layer deposition (ALD with Trimethylalluminum and H2O precursors) using frequency dependent capacitance measurements, where the back side of the samples was coated with a silver paste to form a better contact. The Ti-Pd bilayer has been extensively investigated for gating applications in the context of sensitive electronic devices (e.g., electron and hole spin qubits [46, 47]), and facilitates evaporation and lift-off processes while exhibiting low charge noise. Consequently, it serves as a valuable benchmark for a comparison with the Cr-Co bilayer. An example of C(f)-V measurements for a cobalt MOS with thermal SiO2 is reported in Figure 3a. The effective metal work function () and density of traps (Neff) have been computed by comparing the high frequency characteristics at 1 MHz for different oxide thicknesses () using the method [48] and Equation (1), where the flat-band voltage (VFB) is extracted from the experimental total capacitance Ctot(V measured as a function of the applied bias (VG). Cox is the oxide capacitance measured in accumulation, the doping of the substrate, the elementary charge, the semiconductor potential, and the dielectric constant of silicon/oxide, respectively:
@@ 3 FDSOI Nanowires with Ferromagnetic Cobalt Gates
To demonstrate the co-integration of ferromagnetic gates with standard metallic control gates, an FDSOI nanowire device comprising two quantum dots has been fabricated (Figure 1). The top 100 silicon layer facing the buried-oxide was selectively etched down to 18 nm and the nanowire patterned by e-beam lithography and reactive ion etching with SF6. The width of the nanowire was set to 50 nm to facilitate electrostatic and physical quantum confinement, whereas wider nanowires should be adopted for the control of separated corner dots with split gates, as successfully reported for quantum dots with holes [54]. Palladium plunger gates with a thin adhesion layer of titanium have been fabricated to control the chemical potential of the dots in the nanowire. They alternate with barrier gates made of chromium-cobalt, used to electrostatically tune depletion tunneling barriers for quantum confinement, and are insulated by ALD-Al2O3. Ohmic contacts have been fabricated with phosphorus-implanted n wells () contacted with Ti/Pt to probe electron conduction through the structure from room temperature to mK with minimal parasitic series resistance. To isolate the quantum dots from diffused dopants, two lateral Pd inversion gates extend the highly conductive implanted regions to the dots area, creating two continuous inversion layers alongside the nanowire. All layers have been fabricated with e-beam lithography, metal e-beam evaporation and lift-off process. The average pitch is 70 nm and a barrier gate 60 nm wide alternates with a pair of 35 nm Co gates, to tune the addressability of qubits. This architecture is compatible with a self-aligned dry etching process [55], an effective solution to enable large-scale integration of quantum dots controlled by nanomagnets.
The independent electrical control of the two layers of gates at room temperature is reported in Figure 4 (a-d). Each layer exhibits pinch-off curves with subthreshold swing (SSRT) below 100 mV/dec and ohmic behaviour of the contacts. A threshold voltage mismatch () is measured between the two layers, due to the interlayer oxide and difference in work-function. Finally, the linear threshold voltage dependence of barrier gates on the plunger’s bias (and vice versa) suggests a low inter-layers defectivity [56], validating the quality of insulation between the metals. Figure 4 (e-h) reports similar curves for a plunger gate and two barrier gates defining a quantum dot at cryogenic temperature (10 mK). Here, the discrepancy between the characteristics of the two barriers stems from the varying dimensions of the cobalt gates (which result in a different capacitance to the semiconductor), that were specifically selected to induce varying stray fields in the region below the plunger gates, leading to different Larmor frequencies in the two quantum dots.

### v2

@@ Abstract
The rise of electron spin qubit architectures for quantum computing processors has led to a strong interest in designing and integrating ferromagnets to induce stray magnetic fields for electron dipole spin resonance (EDSR). The integration of nanomagnets imposes however strict layout and processing constraints, challenging the arrangement of different gating layers and the control of neighboring qubit frequencies. This work reports a successful integration of nano-sized cobalt control gates into a multi-gate fully-depleted silicon-on-insulator (FDSOI) nanowire with nanometer-scale dot-to-magnet pitch, simultaneously exploiting electrical and ferromagnetic properties of the gate stack at nanoscale. The electrical characterization of the multi-gate nanowire exhibits full field effect functionality of all ferromagnetic gates from room temperature to 10 mK, proving quantum dot formation when ferromagnets are operated as barrier gates. The front-end-of-line (FEOL) compatible integration of cobalt is examined by energy dispersive X-ray spectroscopy and high/low frequency capacitance characterization, confirming the quality of interfaces and control over material diffusion. Insights into the magnetic properties of thin films and patterned control-gates are provided by vibrating sample magnetometry and electron holography measurements. Micromagnetic simulations anticipate that this structure fulfills the requirements for EDSR driving for magnetic fields higher than 1 T, where a homogeneous magnetization along the hard magnetic axis of the Co gates is expected. The FDSOI architecture showcased in this study provides a scalable alternative to micromagnets deposited in the back-end-of-line (BEOL) and middle-of-line (MOL) processes, while bringing technological insights for the FEOL-compatible integration of Co nanostructures in spin qubit devices.
@@
ferromagnet, cobalt, CMOS integration, quantum dots, EDSR, spin qubitsFabio Bersano* Michele Aldeghi Niccolò Martinolli Victor Boureau Thibault Aboud Michele Ghini Pasquale Scarlino Gian Salis Adrian M. Ionescu
F. Bersano, N. Martinolli, T. Aboud, Dr. M. Ghini, Prof. A. M. Ionescu
Nanoelectronic Devices Laboratory (Nanolab), IEM EPFL, 1015 Lausanne, Switzerland
Email Address: fabio.bersano@epfl.ch, adrian.ionescu@epfl.ch
M. Aldeghi, Dr. G. Salis
IBM Research-Zurich, 8803 Rüschlikon, Switzerland
Dr. Victor Boureau
Interdisciplinary Center for Electron Microscopy, EPFL, 1015 Lausanne, Switzerland
Prof. P. Scarlino
Hybrid Quantum Circuits Laboratory (HQC Lab), IPHYS EPFL, 1015 Lausanne, Switzerland
@@ 1 Introduction
Particularly for electron spin qubits in Si, the need for plunger, depletion/barrier gates and micromagnet integration puts stringent restrictions on layout level when scaling up the quantum processor, increasing the complexity of gating layers for electrostatic control. This challenge can be addressed by properly designing the device architecture and combining the magnetic and electrical functionality of ferromagnetic control gates [28, 29, 30]. While a 2D architecture proves more fitting for expansive networks [31], it necessitates additional depletion gates to induce 1D confinement, a role that is naturally fulfilled by etched structures like FinFETs and nanowires, which typically require a minimum of three metal layers [25, 10].
In this work, we propose a device architecture designed to address the scalability challenges of EDSR-driven electron spin qubits, enhancing both driving gradient and addressability of neighbouring qubits, while minimizing decoherence gradients. In our structure, the 1D confinement of the quantum dots is achieved using ultra-thin fully-depleted silicon-on-insulator (FDSOI) nanowires. Ferromagnetic gates are integrated to produce slanting Zeeman fields while controlling the local depletion of minority carriers within the SOI nanowires through a DC bias, which could be combined with RF excitation (Figure 1a). Two metal bi-layers, Ti/Pd and Cr/Co, act as plunger and barrier gates, respectively. Palladium is chosen for its small grain size [32], which facilitates patterning via lift-off, and for its resistance to self-oxidation, unlike aluminum. Cobalt is selected as the ferromagnetic element due to its compatibility with industrial processes for integrated circuits and higher saturation magnetization compared to nickel. Although Fe and its permalloys exhibit higher magnetic saturation and typically lower magnetocrystalline anisotropy than Co [33, 34], they are excluded due to iron being considered one of the most problematic contaminants in metal-oxide-semiconductor devices [35]. The incorporation of Co nano-gates suggested here is first analyzed with an emphasis on its compatibility with FEOL and middle-of-line (MOL) CMOS processes. The fabrication and electrical characterization of a proof-of-concept quantum dot device (Figure 1b) is reported, and micromagnetic simulations are discussed to anticipate the experimental conditions required for EDSR driving induced by Co barrier gates.
@@ 2.1 Thin Film Morphology and Thermal Budget
In the context of semiconductor devices manufacturing, cobalt is considered a CMOS-compatible material. It has been explored as a replacement for traditional interconnects metals like tungsten and copper due to its low electron mean free path [36], and as a valid alternative to TiSi2 and NiSi to form low-resistivity silicide contacts (CoSi2) in advanced technology nodes [37]. Co interconnects have been successfully integrated for CMOS nodes below 100 nm in the MOL metallization layers and as cladding coating at higher metal levels, showing more than two times reduction in via resistance [38] and 5-10 times improvement in electromigration [39]. However, instances of extrinsic time-dependent dielectric breakdown and atomic diffusion have been documented, prompting the need for diffusion barriers [40, 41]. Even a partial diffusion of Co atoms into the gate oxide and silicon could result in the creation of interfacial silicides [42], which may lead to the development of local magnetic moments with random orientations compared to the magnetization of the thin film. Yet, it has been demonstrated that a thin SiO2 layer effectively inhibits the formation of interfacial CoSi/CoSi2 due to the strong binding of Si atoms to the thermal SiO2 layer [43]. However, aluminum oxide does not exhibit this diffusion barrier effect [40], necessitating the use of an additional adhesion layer to separate the Co from the gate oxide. Because of the partial diffusion of Cr into high-k dielectrics and reaction with oxygen to form Cr2O3, a Cr-Co alloy provides a reliable self-forming diffusion barrier that shows a breakdown voltage two times higher than pure Co [44]. Motivated by this, we first deposit a thin Cr layer (2-3 nm) before Co e-beam evaporation; this layer also enhances metal adhesion to the gate oxide and improves the lift-off process. The cross-section of a Cr/Co barrier gate and Ti/Pd plunger gate annealed for 10 minutes at 300 °C in forming gas (5 H2, 95 N2) are shown in Figure 2a, where energy dispersive X-ray (EDX) spectroscopy in a transmission electron microscope (TEM) was used to highlight the materials. No Cr/Co or Ti/Pd diffusion was observed in the oxide, indicating effective atomic screening by the thin adhesion metal, though partial diffusion of the adhesion layers into Co and Pd occurred.
In addition to inter-layer diffusion, the oxidation of Co must be considered since it occurs rapidly when in contact with oxygen [45]. Three oxide growth processes have been documented: an immediate hydroxide formation (Co(OH)2) of a few Ångstrom ( nm) upon air exposure of the surface at annealing temperatures below 100 °C, a logarithmic growth of CoO for annealing temperatures between 100-225 °C, and a quadratic growth rate of a combination of CoO and Co3O4 for temperatures exceeding 300 °C [46, 47]. We analyzed the oxidation phase of Co through angle-resolved X-ray photoelectron spectroscopy (ARXPS) of continuous thin films (3/33 nm of Cr/Co) deposited on both SiO2 and Al2O3 before and after rapid thermal annealing (RTP). Results are independent of the dielectric used and are reported in Figure 2b for (detection normal to the surface). Upon collection of a high resolution spectrum in the interval typical of cobalt, two sets of peaks separated by eV are measured, identifying the binding energy profile of Co and its oxides in accordance with data reported in literature [48], with a sharp increase in the metallic peak measured after RTP as typical of surface dehydration. The detection at lower binding energies identifies two peaks corresponding to Cr2O3 in the sample subjected to RTP at °C, providing experimental evidence of Cr diffusion into Co and its segregation at the outer interface. This is further confirmed by a quantitative analysis at higher detection angles and EDX-TEM for a nanogate cross-section (details in SI). Vibrating sample magnetometry (VSM) measurements were conducted to study the evolution of coercivity with rapid thermal treatments in forming gas, results are shown in Figure 2c. In agreement with previous studies on thermally induced changes of structural and magnetic properties in Co thin films [49], an abrupt increase in coercive field is measured at an annealing temperature of 400 °C. This behavior was previously attributed to improvements in the crystalline structure and magneto-crystalline anisotropy of the growing fcc crystallites, along with enhanced roughness at the metal-dielectric interface. We studied the evolution of crystallite size and crystal phase of thin Co evaporated films by bright-field (BF) TEM and selected area electron diffraction (SAED) measurements [50, 51]. The results, shown in Figure 2d, indicate a grain size ranging from 15 to 20 nm, which remains mostly unaffected by the thermal treatment at °C, but roughly doubles in size at °C. The higher intensity peaks correspond to the hexagonal close packed (hcp) phase, as confirmed by grazing incidence XRD (GIXRD) measurements. After annealing at 300 °C, an improvement in the crystalline structure and stress relaxation are attained, as suggested by the reduction in diffuse brightness around the main peaks, which in turn highlights the presence of face centered cubic (fcc) phase. The fcc phase appears to slightly increase in the sample annealed at °C, and this, together with the crystallites growth and possibly changes in the interface, could contribute to explain the observed rise in magnetic coercivity. The decreased uniformity in the SAED rings has to be attributed to the smaller number of crystallites framed by the same BF field of view, rather than to any film texturing.
Both metallization layers of the multi-gate architecture proposed in this work are patterned on thin dielectrics deposited via atomic layer deposition (ALD) at °C. We observed a high risk of breaking the nanogates when annealed at temperatures above the dielectric deposition temperature, due to stress generated at the dielectric-metal interface. During all annealing processes conducted up to 300 °C, surface oxidation of the Co nanogates was limited to 4 nm by cooling the sample to room temperature in a nitrogen atmosphere before exposing it to ambient conditions. Our study thus recommends 300 °C as a safe thermal limit for integrating Co nanomagnets on ALD-dielectrics (which are typically deposited at lower temperatures), as it minimizes the risk of magnetic inhomogeneities caused by high oxidation rates and interfacial stress between the ferromagnets and dielectric. Further characterization studies should explore the feasibility to extend the thermal budget to °C (typical for CMOS back-end processing) by patterning Co nanogates on thermally oxidized SOI nanowires.
@@ 2.2 Work Function and Interface Traps
The primary criteria for incorporating a metal as gating layer include its chemical stability with a given dielectric to regulate the threshold voltage (Vth), negligible C-V hysteresis, and low interface traps. All these properties depend on the thermal history of the fabrication process and set boundaries to the thermal budget. We analyzed the density of traps in metal-oxide-semiconductor (MOS) capacitors fabricated with Cr-Co (3-33 nm) and Ti-Pd (3-33 nm) on thermal SiO2 (grown at 850 °C in dichloroethene atmosphere) and Al2O3 deposited by ALD (with trimethylalluminum and H2O precursors) using frequency dependent capacitance measurements, where the back side of the samples was coated with a silver paste to form a better contact. The Ti-Pd bilayer has been extensively investigated for gating applications in the context of sensitive electronic devices (e.g., electron and hole spin qubits [52, 53]), and facilitates evaporation and lift-off processes while exhibiting low charge noise. Consequently, it serves as a valuable benchmark for a comparison with the Cr-Co bilayer. An example of C(f)-V measurements for a cobalt MOS with thermal SiO2 is reported in Figure 3a. The effective metal work function () and density of traps (Neff) have been computed by comparing the high frequency characteristics at 1 MHz for different oxide thicknesses () using the method [54] and Equation (1), where the flat-band voltage (VFB) is extracted from the experimental total capacitance Ctot(V measured as a function of the applied bias (VG). Cox is the oxide capacitance measured in accumulation, the doping of the substrate, the elementary charge, the semiconductor potential, and the dielectric constant of silicon/oxide, respectively:
@@ 3 FDSOI Nanowires with Ferromagnetic Cobalt Gates
To demonstrate the co-integration of ferromagnetic gates with standard metallic control gates, an FDSOI nanowire device comprising two quantum dots has been fabricated (Figure 1). The top 100 silicon layer facing the buried-oxide was selectively etched down to 18 nm and the nanowire patterned by e-beam lithography and reactive ion etching with SF6/C4F8. The width of the nanowire was set to 50 nm to facilitate electrostatic and physical quantum confinement, whereas wider nanowires should be adopted for the control of separated corner dots with split gates, as successfully reported for quantum dots with holes [60]. Palladium plunger gates with a thin adhesion layer of titanium have been fabricated to control the chemical potential of the dots in the nanowire. They alternate with barrier gates made of chromium-cobalt, used to electrostatically tune depletion tunneling barriers for quantum confinement, and are insulated by ALD-Al2O3. Ohmic contacts have been fabricated with phosphorus-implanted n++ wells () contacted with Ti/Pt to probe electron conduction through the structure from room temperature to millikelvin with minimal parasitic series resistance. To isolate the quantum dots from diffused dopants, two lateral Pd inversion gates extend the highly conductive implanted regions to the dots area, creating two continuous inversion layers. All layers have been fabricated with e-beam lithography, metal e-beam evaporation and lift-off process. The average pitch is 70 nm and a barrier gate 60 nm wide alternates with a pair of 35 nm Co gates; this arrangement creates varying stray fields in the region beneath the plunger gates, resulting in different Larmor frequencies in the two quantum dots (see subsection 3.2.1). It is worth mentioning that industrial processing relies on dry etching techniques, and Co can be etched using such methods [61], providing a promising solution for large-scale integration of nanomagnets.
The independent electrical control of the two layers of gates at room temperature is reported in Figure 4 (a-d). Each layer exhibits pinch-off curves with subthreshold swing (SSRT) below 100 mV/dec and ohmic behaviour of the contacts. A threshold voltage mismatch () is measured between the two layers, due to the interlayer oxide and difference in work-function. Finally, the linear threshold voltage dependence of barrier gates on the plunger’s bias (and vice versa) suggests a low inter-layers defectivity [62], validating the quality of insulation between the metals. Figure 4 (e-h) reports similar curves for a plunger gate and two barrier gates defining a quantum dot at cryogenic temperature (10 mK). Here, the discrepancy between the characteristics of the two barriers stems from the varying dimensions of the cobalt gates which result in a different capacitance to the semiconductor.